# Singular-degenerate parabolic systems with the conormal boundary condition on the upper half space

### v1

@@ 6. Equations with singular lower order terms
We now give the proof of Remark 2.4 (6). That is, that from $w_{3}(x_{d})\in Y_{p}(\mu)$ it follows that any interior weak solution $u$ to (1.1) with $u,Du,F,f\in L_{q,p}(\Omega_{T},wd\mu)$ automatically satisfies the conormal boundary condition (1.2). Indeed, fix any $\varphi\in C_{0}^{\infty}((-\infty,T)\times\mathbb{R}^{d})$. Denote by $\tilde{w}(t,x)=w_{0}(t)^{-\frac{1}{q-1}}w_{2}(x^{\prime})^{-\frac{1}{p-1}}w_{3}(x_{d})^{-\frac{1}{p-1}}$ the weight dual to $w$. By Hölder’s inequality, for any $f,g$, it holds that
$\left|\int_{\Omega_{T}}f\cdot gd\mu\right|\leq\|f\|_{L_{q,p}(\Omega_{T},wd\mu)}\|g\|_{L_{q^{\prime},p^{\prime}}(\Omega_{T},\tilde{w}d\mu)}.$
As mentioned in Remark 2.4 (2), $w_{3}(x)^{-\frac{1}{p-1}}\in X_{p^{\prime}}(\mu)$. We have that (2.3) holds with $\varphi_{k}$ instead of $\varphi$, where $\varphi_{k}=\varphi(t,x)\zeta(kx_{d})\in C_{0}^{\infty}(\Omega_{T})$ is the approximating sequence from Lemma 6.3. By Lemma 6.3, $\partial_{t}\varphi_{k},D\varphi_{k},\varphi_{k}/x_{d},\varphi_{k}$ converge in $L_{q^{\prime},p^{\prime}}(\Omega_{T},\tilde{w}d\mu)$ to respectively $\partial_{t}\varphi,D\varphi,\varphi/x_{d},\varphi$. Passing to the limit as $k\to\infty$, identity (2.3) holds for any $\varphi\in C_{0}^{\infty}((-\infty,T)\times\mathbb{R}^{d})$.

### v2

@@ 6. Equations with singular lower order terms
We now give the proof of Remark 2.4 (6). That is, in Theorem 2.3 from $w_{3}(x_{d})\in Y_{p}(\mu)$ it follows that any interior weak solution $u$ to (1.1) with $u,Du,F,f\in L_{q,p}(\Omega_{T},wd\mu)$ automatically satisfies the conormal boundary condition (1.2). Indeed, fix any $\varphi\in C_{0}^{\infty}((-\infty,T)\times\mathbb{R}^{d};\mathbb{R}^{n})$. Denote by $\tilde{w}(t,x)=w_{0}(t)^{-\frac{1}{q-1}}w_{2}(x^{\prime})^{-\frac{1}{p-1}}w_{3}(x_{d})^{-\frac{1}{p-1}}$ the weight dual to $w$. By Hölder’s inequality, for any $f,g$, it holds that
$\left|\int_{\Omega_{T}}f\cdot gd\mu\right|\leq\|f\|_{L_{q,p}(\Omega_{T},wd\mu;\mathbb{R}^{n})}\|g\|_{L_{q^{\prime},p^{\prime}}(\Omega_{T},\tilde{w}d\mu;\mathbb{R}^{n})}.$
As mentioned in Remark 2.4 (2), $w_{3}(x_{d})^{-\frac{1}{p-1}}\in X_{p^{\prime}}(\mu)$. We have that (2.3) holds with $\varphi_{k}$ instead of $\varphi$, where $\varphi_{k}=\varphi(t,x)\zeta(kx_{d})\in C_{0}^{\infty}(\Omega_{T})$ is the approximating sequence from Lemma 6.3. By Lemma 6.3, $\partial_{t}\varphi_{k},D\varphi_{k},\varphi_{k}/x_{d},\varphi_{k}$ converge in $L_{q^{\prime},p^{\prime}}(\Omega_{T},\tilde{w}d\mu)$ to respectively $\partial_{t}\varphi,D\varphi,\varphi/x_{d},\varphi$. Passing to the limit as $k\to\infty$, identity (2.3) holds for any $\varphi\in C_{0}^{\infty}((-\infty,T)\times\mathbb{R}^{d};\mathbb{R}^{n})$.
7. Hilbert space-valued solutions
We finish the paper by discussing the extension of our results to the general case of Hilbert space-valued solutions. Let $H$ be a Hilbert space over $\mathbb{R}$ or $\mathbb{C}$, not necessarily separable. Let $\langle\cdot,\cdot\rangle_{H}$ denote the Hermitian form on $H$, which we assume is antilinear in the second coordinate. Let $\mathcal{B}(H)$ be the algebra of bounded linear operators on $H$. If $\in\mathcal{B}(H), aa^{*}$ stands for the operator adjoint to $a$.
Throughout the paper, all of our main and intermediate results continue to hold almost verbatim with $\mathbb{R}^{n}$ changed to $H$, $\mathbb{R}^{n\times n}$ changed to $\mathcal{B}(H)$, and the following modifications in mind. All dot products $u\cdot v$ should be understood as $\langle u,v\rangle_{H}$. All vector-valued functions $u,D_{i}u,F^{i},F_{0},f:\Omega_{T}\to H$ etc. are strongly measurable in the sense of Bochner and all integrals of vector-valued functions are Bochner integrals. We refer to [30, Chapter 1] for the definitions of Bochner integral and measurability for vector-valued functions.
We now discuss the measurability conditions on the operator-valued coefficients. Here a distinction between $H$ being separable or non-separable needs to be made. If $H$ is non-separable, we assume that $a_{0},A^{ij},B^{i},\hat{B}^{i},C,C_{0}:\Omega_{T}\to\mathcal{B}(H)$ etc. are strongly measurable with respect to the norm topology on $\mathcal{B}(H)$ (i.e., in the sense of Bochner), which is a usual assumption in the theory of parabolic PDEs in Banach spaces. If $H$ is separable, we impose a weaker assumption that $a_{0},A^{ij},B^{i},\hat{B}^{i},C,C_{0}:\Omega_{T}\to\mathcal{B}(H)$ etc. are strongly measurable with respect to the strong operator topology. That is, for every $v\in H$ the $H$-valued function $C_{0}v:\Omega_{T}\to H$ is strongly measurable, and similarly for the other coefficients. We impose these conditions in order for the scalar functions $|a_{0}|,|A|,|B|,|\hat{B}|,|C|,|C_{0}|$ to be measurable on $\Omega_{T}$. If $\operatorname{dim}H<\infty$, then $\operatorname{dim}\mathcal{B}(H)<\infty$ and therefore both conditions are equivalent to the usual Lebesgue measurability.
Next, all matrix transposes such as $a_{0}^{T}$ are understood as $a_{0}^{*}$. If the base field is $\mathbb{C}$, we modify (2.1) and (2.2) as follows:
$\kappa\sum_{i=1}^{d}|\xi|_{H}^{2}\leq\operatorname{Re}\sum_{i,j=1}^{d}\langle A^{ij}\xi_{j},\xi_{i}\rangle_{H},\quad\kappa\sum_{i=1}^{d}|\xi|_{H}^{2}\leq\operatorname{Re}\langle C_{0}\xi,\xi\rangle_{H}.$
In the proofs of Theorems 3.1 and 3.2 in Section 3, we change all integrals to their real parts. Starting from Section 4, the proofs carry forward verbatim by changing $\mathbb{R}^{n}$ to $H$ and with the above conventions in mind.